Why do economists believe that inflation in the long run is entirely a monetary phenomenon? Explain fully.
Answer #1

Under the monetary phenomenon, velocity of money is considered as constant.

Then, as per the quantity theory of money in monetarism,

M*V = P*Y = Nominal GDP

So, when V is fixed and M as money supply increases, then nominal GDP also increases. Though, it does not explain whether it is due to the increase in output or increase in the price only. So, inflation will take place when increase in money supply will outpace the increase in output level in the economy. It is considered as monetary phenomenon and it is recognized by the economists in the long run. Further, price is a nominal variable that can only be changed in the long run as output in the long run is fixed that is output at full employment. Hence, increase in money supply leads economists to say that inflation is a monetary phenomenon in the long run.